# REArtGS++: Generalizable Articulation Reconstruction with Temporal Geometry Constraint via Planar Gaussian Splatting

### v2

@@ 5.1 Experimental Setting
Datasets. We compare our method with existing SOTAs on three datasets: (1) PARIS dataset [jiayi2023paris], including 10 objects from PartNet-Mobility dataset [SAPIEN]; (2) ArtGS-Multi [liu2025building], including 5 objects from PartNet-Mobility dataset, and We collect 2 additional screw joint objects from PartNet-Mobility dataset; (3) Real-world dataset, containing 2 objects from PARIS and 3 objects collected by ours. Each state provides 60-100 posed RGB images.
Metrics. Following ArtGS, we generate meshes of both start and end states from canonical state, and report the mean Chamfer Distance (CD) resuls, which are multiplied by 1,000. We employ CD-w, CD-s, and CD-m for whole surfaces, static parts and dynamic parts respectively. We evaluate the predicted joints using angular error (Axis Ang.) and pivot distance (Axis Pos., except for translation joints), which are measured by cm, i.e., multiplied by 100. We also report part motion error, which quantifies the geodesic rotation error (in degrees) and the Euclidean translation distance error (in meters). Lower ($\downarrow$) is better for all metrics.
@@ 5.3 Experiments on Screw and Multi-Part Articulated Objects
Table 3: Ablation of key componets. We report the average results on ArtGS-Multi and 2 screw-joint objects. Axis Pos results are measured by cm. “GS” and “dis” denote Gaussians and distance respectively.
@@ 5.4 Ablation Study
Ablation of key components. To further validate the effectiveness of our key components, we conducted ablation studies on ArtGS-Multi dataset with 2 additional objects exhibiting screw joints. We present the quantitative results in Table. 3 and a qualitative comparison of local voting regularization in Fig. 6. For “w/o planar GS”, we employ the original 3DGS pipeline [kerbl3Dgaussians]; for “w/o screw motion”, joint motions are represented with dual quaternions; for “w/o initialization”, all part centers and joint pivots are randomly initialized. Experimental results demonstrate that each component contributes significantly to overall performance.
@@ 6 Conclusion and Limitations
Our main limitations lie in transparent object reconstruction and the requirement for camera pose alignment between two states. Future work will combine depth estimation and transparent material modeling methods to address transparent surface perception [depthanything, li2025tsgs], and jointly refine two-state camera poses [hong2024pf3plat] in real world data.

### v3

@@ 5.1 Experimental Setting
Datasets. We compare our method with existing SOTAs on three datasets: (1) PARIS dataset [12], including 10 objects from PartNet-Mobility dataset [25]; (2) ArtGS-Multi [13], including 5 objects from PartNet-Mobility dataset, and We collect 2 additional screw joint objects from PartNet-Mobility dataset; (3) Real-world dataset, containing 2 objects from PARIS and 3 objects collected by ours. Each state provides 60-100 posed RGB images.
Metrics. Following ArtGS, we generate meshes of both start and end states from canonical state, and report the mean Chamfer Distance (CD) resuls, which are multiplied by 1,000. We employ CD-w, CD-s, and CD-m for whole surfaces, static parts and dynamic parts respectively. We evaluate the predicted joints using angular error (Axis Ang.) and pivot distance (Axis Pos., except for translation joints), which are measured by mm, i.e., multiplied by 100. We also report part motion error, which quantifies the geodesic rotation error (in degrees) and the Euclidean translation distance error (in meters). Lower ($\downarrow$) is better for all metrics.
@@ 5.3 Experiments on Screw and Multi-Part Articulated Objects
Table 3: Ablation of key componets. We report the average results on ArtGS-Multi and 2 screw-joint objects. Axis Pos results are measured by $m$. “GS” and “dis” denote Gaussians and distance respectively.
@@ 5.4 Ablation Study
Ablation of key components. To further validate the effectiveness of our key components, we conducted ablation studies on ArtGS-Multi dataset with 2 additional objects exhibiting screw joints. We present the quantitative results in Table. 3 and a qualitative comparison of local voting regularization in Fig. 6. For “w/o planar GS”, we employ the original 3DGS pipeline [9]; for “w/o screw motion”, joint motions are represented with dual quaternions; for “w/o initialization”, all part centers and joint pivots are randomly initialized. Experimental results demonstrate that each component contributes significantly to overall performance.
@@ 6 Conclusion and Limitations
Our main limitations lie in transparent object reconstruction and the requirement for camera pose alignment between two states. Future work will combine depth estimation and transparent material modeling methods to address transparent surface perception [26, 11], and jointly refine two-state camera poses [4] in real world data.